OpenStudy (ilovemath14):

HELP ASAP!! FAN AND MEDAL!! 1. x^2+x-12 2. x^2+5x-36= (x-4) ( [ ] ) 3. x^2-12x-28 4. x^2+2x-8 5. x^2-4x-45

OpenStudy (ilovemath14):

@MissSmartiez @mathmale @SnuggieLad

OpenStudy (jdoe0001):

oooooooooooooooooook

OpenStudy (ilovemath14):

What? @jdoe0001

OpenStudy (ilovemath14):

@SkylyerBailey

OpenStudy (skylyerbailey):

I don't know this, I am not good at math lol, sorry

OpenStudy (ilovemath14):

its ok @SkylyerBailey

OpenStudy (ilovemath14):

@Trisarahtops @truwhovian11 @Mathsman @mathexpert

OpenStudy (misssmartiez):

@rebeccaxhawaii

OpenStudy (misssmartiez):

@retirEEd

rebeccaxhawaii (rebeccaxhawaii):

whats the question

OpenStudy (jdoe0001):

well, you're not asking anything, you simply posted a few polynomials

rebeccaxhawaii (rebeccaxhawaii):

^^^^

OpenStudy (ilovemath14):

Factoring Trinomials @rebeccaxhawaii @jdoe0001

jhonyy9 (jhonyy9):

x^2 +x -12 = ? how you factoriz ? do you know the way ?

OpenStudy (fortytherapper):

Lets do number one, then you can use that model for the rest. All of these are quadratics, which means the answer will be (X + or - A)(X + or - B) So for X^2 + X - 12, what do you think is the way to find A and B?

jhonyy9 (jhonyy9):

x^2 +Sx -P = where S is the sum of two numbers and P is the product of these numbers can you getting these numbers ?

jhonyy9 (jhonyy9):

you can writing a system of equation a+b= 1 a*b=-12 can you solve it ?

jhonyy9 (jhonyy9):

or what numbers are the factors of 12 what assumed result 1 ?

jhonyy9 (jhonyy9):

sorry of - 12

OpenStudy (ilovemath14):

3. x^2-12x-28 / \ x^2-14x+2x-28 x(x-14) + 2(x-14)

OpenStudy (ilovemath14):

@jhonyy9

jhonyy9 (jhonyy9):

so this is right than the first two how will be ?

OpenStudy (ilovemath14):

(x+2) (x-14) @jhonyy9

jhonyy9 (jhonyy9):

yes right

jhonyy9 (jhonyy9):

so this is easy - yes ?

OpenStudy (ilovemath14):

ok thank you @jhonyy9

jhonyy9 (jhonyy9):

was my pleasure good luck bye and anytime than i m on line